steve816:

Help solve for x 33x = 60 (mod 97)

moreso:

x = 90

steve816:

I know that's the answer, but how did you get that?

steve816:

I found thatt 33 inverse of mod 97 is 50. What do I do from there?

myininaya:

33x-60=97k k is an integer 33x-97k=60 97=33(2)+31 33=31(1)+2 so 33-31(1)=2 (rewrote last equation above) 33-[97-33(2)]=2 (rewrote next to last above) Combining like terms: 3(33)-97=2 60 is a multiply of 2 2 times 30 is 60 multiply both sides by 30 90(33)-30(97)=60 Does that help? I used euclidean algorithm. Can you see what x is?

steve816:

Ohh yes, I get it now. Thanks for the help @myininaya